OpenStudy (anonymous):

For a recent project, a teacher purchased 250 pieces of red construction paper and 114 pieces of blue construction paper. What is the ratio of red to blue? a. 57:125 c. 125:182 b. 125:57 d. 182:125

OpenStudy (mayankdevnani):

\[\huge \frac{\color{red}{red}}{\color{blue}{blue}}\] \[\huge \frac{\color{red}{250}}{\color{blue}{114}}\] \[\huge \frac{\color{red}{125}}{\color{blue}{57}}\]

OpenStudy (mayankdevnani):

so option c is correct answer